What speed should a galaxy move with respect to us so that the sodium line at 589.0 nm is observed at 589.6 nm?

Grade:11

1 Answers

Chandra
6 years ago
Change in wavelength ∆=589.6-589=0.6nmVelocity V=(change in wavelength/original wavelength)Vel. Of light.Velocity V =(0.6÷589)3*10^8=306 km/s